# Repeated Padding as Data Augmentation for Sequential Recommendation

### v1

@@ 6. Conclusion
In this work, we start from the commonly used padding strategy, aiming to fully utilize the idle input space to improve the model performance and training efficiency further. Based on this idea, we propose RepPad, which leverages the original sequence as padding content instead of the special value 0. The RepPad can be considered as a sequence-level data augmentation method. Compared with existing methods, our method does not contain any parameters or training processes, and it is a data augmentation plugin with high applicability. Extensive experiments demonstrate the superiority of RepPad. It improves the performance of various sequential models and outperforms existing data enhancement methods. We also provide multiple perspectives and insights into what makes RepPad effective. For future work, we hope to provide theoretical underpinnings for RepPad or explore more appropriate padding strategies. We are also interested in combining RepPad with other data augmentation methods.

### v2

@@ 6. Conclusion
In this work, we start from the commonly used padding strategy, aiming to utilize the idle input space to further improve the model performance and training efficiency. Based on this idea, we propose RepPad, which leverages the original sequence as padding content instead of the special value 0. The RepPad can be considered as a sequence-level data augmentation method. Compared with existing methods, our method does not contain any parameters or training processes, and it is a data augmentation plugin with high applicability. Extensive experiments demonstrate the superiority of RepPad. It improves the performance of various sequential models and outperforms existing data enhancement methods. We also provide multiple perspectives and insights into what makes RepPad effective. For future work, we hope to provide theoretical underpinnings for RepPad or explore more appropriate padding strategies. We are also interested in combining RepPad with other data augmentation methods.
Acknowledgements.
This work is partially supported by the National Natural Science Foundation of China under Grant No. 62032013, the Science and technology projects in Liaoning Province (No. 2023JH3/10200005), and the Fundamental Research Funds for the Central Universities under Grant No. N2317002.